OpenStudy (anonymous):

slope intercept form of a line parellel to y=2x-7 passing through (-2,1)

OpenStudy (anonymous):

wats the question?

OpenStudy (anonymous):

find the slope intercept form of a line parallel to y=2x-7 and passing through (-2,1)

OpenStudy (anonymous):

can anyone help me

OpenStudy (anonymous):

slope intercept form is y=mx+b

OpenStudy (anonymous):

a parallel line will have the same slope of 2. insert the x & y into the equation

OpenStudy (anonymous):

I need the parallel line to the line

OpenStudy (anonymous):

1=2(-2)+b 1=-4+b solve for b 5=b now plug the slope and b into the slope intercept form of the equation

OpenStudy (anonymous):

So we know that that the slop of the new equation has to be the same slope as your original equation which is "2" plug your new slope and the value of the points in this equation: (y+b) = m(x+a) where "a" and "b" is where your points have to pass through (-2,1) and your m = 2

OpenStudy (anonymous):

slope of the given line=2 eq. of any line through (-2,1) and slope m is y-1=m(x+2) here m=2 solve and write in the form y=mx+b

OpenStudy (anonymous):

so it would be (0,5)?

OpenStudy (anonymous):

That's the y-intercept. The equation is: y=2x+5

OpenStudy (anonymous):

so the parallel line would be up 2 over 5?

OpenStudy (anonymous):

you start at (0,5) and move up 2, right 1 By extending the line, it should pass through the given point (1,-2) and be parallel to the given line.

OpenStudy (anonymous):

thank you so much

OpenStudy (anonymous):

You're welcome

OpenStudy (anonymous):

I have another couple questions can you still help me

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

how would I graph that

OpenStudy (anonymous):

my graph only goes to 5

OpenStudy (anonymous):

point on (0,5)--the y intercept. Then go up 2 and right 1. Graphing at least 3 points helps. Then extend the line

OpenStudy (anonymous):

Keep plotting points

OpenStudy (anonymous):

ok

OpenStudy (anonymous):